OpenStudy (anonymous):

solve the systems of equations algebraically... y=4x+9 y=x^2+5x+3

OpenStudy (anonymous):

-4 x+y-9 = 0

OpenStudy (anonymous):

y = x (x+5)+3

OpenStudy (anonymous):

\[\left\{4 x+9=x^2+5 x+3\right\} \]Solve the above for x. x = -3, x = 2

OpenStudy (anonymous):

thanks!